OpenStudy (anonymous):

Factor over the complex numbers, if possible: 12m^2 + 27 A) 3(2m + 3i)^2 B) 3(2m - 3i)^2 C) 3(2m + 3i)(2m -3i) D) not possible

Nnesha (nnesha):

first find the GCF

OpenStudy (anonymous):

\[\begin{align*} 12m^2+27&=3(4m^2+9)\\ &=3((2m)^2-(-9))\\ &=3((2m)^2-(3i)^2) \end{align*}\] Now factor as a difference of squares.